# Entropy-Based Doppler Centroid Estimation and Speckle Noise Reduction for Spaceborne SAR Imaging

### v5

@@ Abstract
In this paper, we present strip-map mode spaceborne Synthetic Aperture Radar (SAR) imaging with the focus on Doppler centroid frequency estimation. The non-zero Doppler centroid frequency is the result of non-zero squint angle which if it is not compensated it can de-focus the image. We present an efficient method based on the entropy of the reconstructed image to estimate the fractional part of the Doppler centroid frequency. Furthermore, we discuss the speckle noise, which degrades the quality of the reconstructed images considerably, and attempt to alleviate its effect efficiently. Following the implementation of the speckle noise reduction algorithm, a significant improvement in the quality of the reconstructed images is achieved.
Finally, we utilize the experimental data gathered from the RADARSAT-1 satellite of Vancouver Canada to verify the accuracy and effectiveness of the proposed techniques.
@@ I Introduction
Synthetic Aperture Radar (SAR) imaging is a unique method to create the effect of a large real aperture synthetically. The high resolution in the azimuth direction in SAR imaging stems from the relative motion between the radar and the target [1, 2, 3, 4, 5]. In this paper, we specifically consider the strip-map mode spaceborne SAR imaging in which the angle between the main axis of the antenna and the terrain to be imaged, remains fixed [3, 2, 6]. Similar to all high resolution radars, in SAR imagery systems, the fine resolution in the range direction is achieved by utilizing the idea of pulse compression. A Linear Frequency Modulated (LFM) signal, also known as chirp signal, is transmitted and the reflected signal is processed using matched filtering technique and as a result a high resolution profile in the range direction is produced [3, 2, 7, 8]. The range resolution depends on the bandwidth of the chirp signal. One of the features of the strip-map mode SAR imaging is the squint angle of the antenna which shifts the spectrum of the signal in the azimuth direction. For this reason, the signal in the azimuth direction is not at base-band. In other words, the average value of the azimuth frequency, which is called Doppler centroid frequency, is a non-zeros value. Non-zero squint angle is the main reason for Range Cell Migration (RCM) phenomenon which if it is not compensated, it degrades the image quality in both range and azimuth directions considerably [3]. The best possible way to estimate the Doppler centroid frequency is by using the received data. Estimating the Doppler centroid frequency by analyzing the geometry does not give rise to an accurate result since the location of the satellite can not be estimated with high precision [3]. Since the sampling rate in the azimuth direction is equal to the Pulse Repetition Frequency (PRF), therefore, if the Doppler centroid frequency exceeds the PRF, there will be ambiguity in its estimation. In fact there will be two parts for the Doppler centroid, namely, fractional part and the integer part.
@@ II Model Description and Image Formation
$s(t,\eta)=\sigma\;w_{r}(t-\frac{2R(\eta)}{c})w_{a}(\eta-\eta_{c})$
$e^{\displaystyle-j4\pi f_{c}\frac{R(\eta)}{c}+j\pi\beta t^{2}+j4\pi\beta t\frac{R(\eta)}{c}},$ (1)
@@
where $p_{r}(x)=\frac{sin(\pi x)}{\pi x}$. Next, we should compensate for the Range Cell Migration (RCM) phenomenon which is the result of coupling between the range and azimuth directions. Consequently, after RCM compensation, we can describe the signal in the Doppler domain as [3]
$S_{\rm rcmc}(t,f_{\eta})$ $=\sigma\;p_{r}(t-\frac{2R_{0}}{c})W_{a}(f_{\eta}-f_{\eta_{c}})\times$ (3)
$e^{\displaystyle-j\frac{4\pi f_{c}R_{0}}{c}}e^{\displaystyle j\pi\frac{f^{2}_{\eta}}{K_{a}}},$
@@ IV Speckle Noise
Center frequency ($\rm GHz$) $\rm f_{c}$ $5.3$
Radar sampling rate ($\rm MHz$) $\rm Fr$ $32.317$
Pulse repetition frequency ($\rm Hz$) $\rm PRF$ $1256.98$
Slant range of first radar sample ($\rm km$) $R_{0}$ $988.65$
FM rate of radar pulse ($\rm MHz/\mu s$) $\rm\beta$ $0.72135$
Chirp duration ($\rm\mu s$) $\rm T$ $41.75$
Satellite velocity ($\rm m/s$) $\rm v$ $7062$
Bandwidth ($\rm MHz$) $\rm b$ $30.116$
@@
Refer to captionAlong-Track [Samples]Slant-Range [Samples](a)Refer to captionAlong-Track [Samples]Slant-Range [Samples](b)
@@
As we mentioned before, if the Doppler centroid frequency is larger than the PRF of the RADAR, there will be an ambiguity in the Doppler centroid frequency estimation. One way to estimate the unambiguous value for the Doppler centroid frequency is by analyzing the trajectory of the strong isolated targets. Fig. 2-(b) shows a blow-up part of the range compressed image shown in Fig. 2-(a) related to a few stationary ships in water. The skew in the trajectory of these targets is due to the non-zero Doppler centroid frequency. The energy of each of these targets is spread over several different range cells. The slope can easily be calculated and is equal to $0.034$ range samples per azimuth samples. To estimate the Doppler centroid frequency we should first multiply the slope by $\rm\frac{c}{2Fr}$ to convert the slope from range samples to range distance and then multiply it by $\rm PRF$ to convert it from azimuth samples to azimuth time. Hence, we have $\frac{dR(\eta)}{d\eta}=198.23\;\rm m/s$. As a result, from (7) the Doppler centroid frequency can be calculated as $f_{\rm dc}=-7009\;\rm Hz$. Fig. 3-(a) illustrates the result of the RCM compensation process. As can be seen from Fig. 3-(a), the energy of the targets have been localized in their corresponding range cells.
@@
Next, we estimate the fractional part of the Doppler centroid frequency which is $f^{\prime}_{\rm dc}$. The fractional part is essential in focusing the energy of the targets in the azimuth direction. Fig. 3-(b) illustrates the power spectrum of the data in azimuth direction versus slow time frequency $f_{\eta}$. In order to reduce the effect of the noise, we have averaged the power spectrum corresponding to $230$ range cells. In order to be able to calculate the frequency component at which the signal reaches its maximum value, we have performed curve fitting. From Fig. 3-(b), we can easily estimate the fractional part of the Doppler centroid frequency as $f^{\prime}_{\rm dc}=520\;\rm Hz$. Finally, at the last stage, we perform the azimuth localization based on (4). Consequently, the reconstructed image based on the Range-Doppler algorithm is obtained which has been presented in Fig. 4.
@@
Refer to captionNormalized EntropyFrequency [Hz](b)Refer to captionNormalized EntropyFrequency [Hz]

### v6

@@ Abstract
In this paper, we present strip-map mode spaceborne Synthetic Aperture Radar (SAR) imaging with the focus on Doppler centroid frequency estimation. The non-zero Doppler centroid frequency is the result of non-zero squint angle which if it is not compensated it can de-focus the image. We present an efficient method based on the entropy of the reconstructed image to estimate the fractional part of the Doppler centroid frequency. Furthermore, we discuss the speckle noise, which degrades the quality of the reconstructed images considerably, and attempt to alleviate its effect efficiently. Following the implementation of the speckle noise reduction algorithm, a significant improvement in the quality of the reconstructed images is achieved. Finally, we utilize the experimental data gathered from the RADARSAT-1 satellite of Vancouver Canada to verify the accuracy and effectiveness of the proposed techniques.
@@
This paper has been published in the 2025 International Conference on Communication and Signal Processing (ICCSP) ©IEEE
@@ I Introduction
Synthetic Aperture Radar (SAR) imaging is a unique method to create the effect of a large real aperture synthetically. The high resolution in the azimuth direction in SAR imaging stems from the relative motion between the radar and the target [1, 2, 3, 4]. In this paper, we specifically consider the strip-map mode spaceborne SAR imaging in which the angle between the main axis of the antenna and the terrain to be imaged, remains fixed [3, 2, 5]. Similar to all high resolution radars, in SAR imagery systems, the fine resolution in the range direction is achieved by utilizing the idea of pulse compression. A Linear Frequency Modulated (LFM) signal, also known as chirp signal, is transmitted and the reflected signal is processed using matched filtering technique and as a result a high resolution profile in the range direction is produced [3, 2, 6, 7]. The range resolution depends on the bandwidth of the chirp signal. One of the features of the strip-map mode SAR imaging is the squint angle of the antenna which shifts the spectrum of the signal in the azimuth direction. For this reason, the signal in the azimuth direction is not at base-band. In other words, the average value of the azimuth frequency, which is called Doppler centroid frequency, is a non-zeros value. Non-zero squint angle is the main reason for Range Cell Migration (RCM) phenomenon which if it is not compensated, it degrades the image quality in both range and azimuth directions considerably [3]. The best possible way to estimate the Doppler centroid frequency is by using the received data. Estimating the Doppler centroid frequency by analyzing the geometry does not give rise to an accurate result since the location of the satellite can not be estimated with high precision [3]. Since the sampling rate in the azimuth direction is equal to the Pulse Repetition Frequency (PRF), therefore, if the Doppler centroid frequency exceeds the PRF, there will be ambiguity in its estimation. In fact there will be two parts for the Doppler centroid, namely, fractional part and the integer part.
@@ II Model Description and Image Formation
$s(t,\eta)=\sigma\;w_{r}(t-\frac{2R(\eta)}{c})w_{a}(\eta-\eta_{c})\times e^{\displaystyle-j4\pi f_{c}\frac{R(\eta)}{c}+j\pi\beta t^{2}+j4\pi\beta t\frac{R(\eta)}{c}},$ (1)
@@
where $p_{r}(x)=\frac{sin(\pi x)}{\pi x}$. Next, we should compensate for the RCM phenomenon which is the result of coupling between the range and azimuth directions. Consequently, after RCM compensation, we can describe the signal in the Doppler domain as [3]
$S_{\rm rcmc}(t,f_{\eta})=\sigma\;p_{r}(t-\frac{2R_{0}}{c})W_{a}(f_{\eta}-f_{\eta_{c}})\times e^{\displaystyle-j\frac{4\pi f_{c}R_{0}}{c}}e^{\displaystyle j\pi\frac{f^{2}_{\eta}}{K_{a}}},$ (3)
@@ V Experimental Results
Center frequency ($\rm GHz$) $\rm f_{c}$ 5.35.35.3
Radar sampling rate ($\rm MHz$) $\rm Fr$ 32.31732.31732.317
Pulse repetition frequency ($\rm Hz$) $\rm PRF$ 1256.981256.981256.98
Slant range of first radar sample ($\rm km$) $R_{0}$ 988.65988.65988.65
FM rate of radar pulse (${\rm MHz}/\mu s$) $\beta$ 0.721350.721350.72135
Chirp duration ($\mu s$) $\rm T$ 41.7541.7541.75
Satellite velocity ($\rm m/s$) $\rm v$ 706270627062
Bandwidth ($\rm MHz$) $\rm b$ 30.11630.11630.116
@@
Refer to captionAlong-Track [Samples]Slant-Range [Samples](a)
Refer to captionAlong-Track [Samples]Slant-Range [Samples](b)
@@
As we mentioned before, if the Doppler centroid frequency is larger than the PRF of the RADAR, there will be an ambiguity in the Doppler centroid frequency estimation. One way to estimate the unambiguous value for the Doppler centroid frequency is by analyzing the trajectory of the strong isolated targets. Fig. 2-(b) shows a blow-up part of the range compressed image shown in Fig. 2-(a) related to a few stationary ships in water. The skew in the trajectory of these targets is due to the non-zero Doppler centroid frequency. The energy of each of these targets is spread over several different range cells. The slope can easily be calculated and is equal to 0.0340.0340.034 range samples per azimuth samples. To estimate the Doppler centroid frequency we should first multiply the slope by $\rm\frac{c}{2Fr}$ to convert the slope from range samples to range distance and then multiply it by $\rm PRF$ to convert it from azimuth samples to azimuth time. Hence, we have $\frac{dR(\eta)}{d\eta}=198.23\;\rm m/s$. As a result, from (7) the Doppler centroid frequency can be calculated as $f_{\rm dc}=-7009\;\rm Hz$. Fig. 3-(a) illustrates the result of the RCM compensation process. As can be seen from Fig. 3-(a), the energy of the targets have been localized in their corresponding range cells.
@@
Next, we estimate the fractional part of the Doppler centroid frequency which is $f^{\prime}_{\rm dc}$. The fractional part is essential in focusing the energy of the targets in the azimuth direction. Fig. 3-(b) illustrates the power spectrum of the data in azimuth direction versus slow time frequency $f_{\eta}$. In order to reduce the effect of the noise, we have averaged the power spectrum corresponding to 230230230 range cells. In order to be able to calculate the frequency component at which the signal reaches its maximum value, we have performed curve fitting. From Fig. 3-(b), we can easily estimate the fractional part of the Doppler centroid frequency as $f^{\prime}_{\rm dc}=520\;\rm Hz$. Finally, at the last stage, we perform the azimuth localization based on (4). Consequently, the reconstructed image based on the Range-Doppler algorithm is obtained which has been presented in Fig. 4.
@@
Refer to captionNormalized EntropyFrequency [Hz](b)
Refer to captionNormalized EntropyFrequency [Hz]
@@ VI Conclusion
As future work, we will focus on comparing the existing methods with the proposed technique for the Doppler centroid frequency estimation and measure the improvement with numeric metrics.